OpenStudy (anonymous):

Describe the usefulness of the conjugate and its effect on other complex numbers. :)

OpenStudy (anonymous):

Well, let's give an example. Here's a complex number: 3+2i To get the complex conjugate of that number, we just change the sign of the imaginary part. 3-2i Now, what is cool about the complex conjugate? Well, let's see what happens when we multiply those two things. (3+2i)(3-2i) FOIL it out. 9 -6i +6i -4i^2 Hey, sweet! The two middle terms cancel out, so I'm left with 9-4i^2 and what do we know about i^2? Well, i^2=-1, so I can substitute and get 9-4(-1) = 9+4 = 13 Wait, what just happened? Well, we multiplied two complex numbers together, and since the middle terms canceled, we were left with just a normal non-complex number.

OpenStudy (anonymous):

That's the cool thing about the complex conjugate. Whenever I multiply a complex number by its conjugate, the result is a real number.